In a certain country license plate numbers have 3 letters followed by 4 digits. How many different license plate numbers can be formed


1
Expert's answer
2020-11-05T16:17:28-0500
"Solution"

There are "3" spots where there are "26" choices each "(=26^3)" and "4" spots where there are "10" choices each (the digits through "9" , "=10^4" ). This gives:


There are "10^4\times26^3 = 175,760,000" different plates possible.